OpenStudy (anonymous):

A cannonball is fired toward a vertical building 480 away with an initial velocity of 120 at 36.9above the horizontal. How many seconds until the cannonball hits the building?

sam (.sam.):



sam (.sam.):

You can use v=d/t here for horizontal since they've given you distance. Find the horizontal velocity. \[v_x=vcos(36.9) \\ \\ v_x=95.96 ms^{-1}\] \[v_x=\frac{d}{t} \\ \\ t=\frac{d}{v_x} \\ \\ t=\frac{480}{95.96} \\ \\ t=5~ seconds\]